# Tiling Spaces and the Expanding Universe: Bridging Quantum Mechanics and Cosmology

### v1

@@ 3 The Universe Expansion from Quasicrystalline Tiling Spaces
The projection of the initial shells of the E8 lattice, when visualized from a 3D slice perspective, reveals patterns that emerge under rescaling. See Image 1 for a visual representation.
@@
The notion of tiling spaces emerges from considering each possible tiling that can be built on a specific CPS as a point in an abstract space. In a quantum mechanical framework, each of these possible tilings can be defined as a state. However, rather than working with individual states, we employ the more comprehensive concept of Hilbert spaces, which encompass all possible states. This approach bridges the geometric nature of tilings with the probabilistic nature of quantum mechanics. Consequently, tiling spaces provide a robust framework for describing both the geometric structure and the quantum mechanical evolution of the universe during its expansion, offering a unified perspective on cosmic dynamics.
@@ 3.1 Universe as a Growing Quasicrystal Hypothesis
This model proposes an unconventional origin of the universe, diverging from the traditional Big Bang theory. It suggests that the universe began with the projection of a single point from a higher-dimensional lattice, such as the E8 root lattice. This initiated a sequence of point projections that sequentially constructed the universe’s structure. This process, governed by a computational framework, unfolds step by step transitioning from one shell $N_{\text{E8}}(n)$ to the next $N_{\text{E8}}(n+1)$ and so forth. The number of points in each shell grows rapidly, with the first ten shells containing 240, 2160, 6720, 17520, 30240, 60480, 82560, 140400, 181680, 272160 points, respectively. The projection into a 4D quasicrystalline structure not only expands the universe but also intricately subdivides it internally, as points from newer shells intermingle with those from older shells, showcasing a complex process of growth and subdivision characteristic of the universe’s early stages (see Figure 2 for a lower-dimensional illustration).
Figure 2: A rotated Z2 being projected shell by shell with the color tracking the ordering. As this one dimensional universe expands, it also get new points inside the already projected region.
In this early stage, the universe consists of points, which are projections from the higher-dimensional root lattice. In the Lie algebra framework of modern gauge theories, these points represent fundamental particle states that form the Hilbert spaces of representations of the Lie algebra. At this stage, there is no scale or mass, and thus no distinction between particles, highlighting an analogy to the grand unification scale.
@@
This growing Hilbert space structure suggests that the early universe may be viewed as a form of a quasicrystalline topological quantum computer, with quantum error-correcting codes (QECCs) inherently embedded [24]. At each growth step, a subspace $\mathcal{C}_{n}\subset\mathcal{H}_{n}$ can be identified, spanned by a set of wavefunctions ${|\Psi{n,[T]}\rangle}$ that exhibit local indistinguishability and recoverability properties [24]:
$|\Psi_{n,[T]}\rangle=\sum_{g\in G_{n}}\beta(gT_{n})|gT_{n}\rangle,$ (24)
This growing QECC structure provides a framework for understanding how quantum information can be encoded and protected within the geometry of a quasicrystalline universe. Notably, the Hilbert spaces associated with certain quasicrystals are isomorphic to the fusion Hilbert spaces that appear in topological quantum computing systems [25]. This establishes a connection between QECCs and topological quantum computation within the context of our working hypothesis of the growing quasicrystalline universe.
The emergence of this quasicrystalline structure and its associated Hilbert spaces in the early universe may have profound implications for the evolution of spacetime, matter, and quantum information processing. As the universe continues to grow and evolve, the interplay between the quasicrystalline geometry, QECCs, and topological quantum computation may guide the formation of large-scale cosmic structures and the emergence of complex phenomena. This heuristic model offers a novel perspective on the origin and evolution of the universe, connecting ideas from condensed matter physics, quantum information theory, and cosmology.
As the quasicrystalline universe evolves, phonons and phasons should emerge naturally due to the intrinsic properties of growing quasicrystals, even those composed of ’spacetime atoms.’ Phonons correspond to collective vibrational modes of the tiling configurations, while phasons are associated with the rearrangement of the quasicrystalline structure without changing the overall energy. The presence of these quasiparticles is significant, as their dispersion relations lead to a form of the Friedmann equation (Equation 13), implying the establishment of standard expansion.
To incorporate phonons and phasons into our quantum description of tiling spaces, we introduce creation and annihilation operators acting on the Hilbert space $\mathcal{H}_{n}$ at each growth step. Let $a^{\dagger}_{k,n}$ and $a_{k,n}$ denote the creation and annihilation operators for a phonon with wave vector $k$ at the $n$-th growth step, satisfying the commutation relations:
@@
The emergence of phonons and phasons in the growing quasicrystalline universe has important implications for the properties of the corresponding quantum error-correcting codes. The presence of these excitations introduces additional degrees of freedom that may affect the error correction capabilities of the code. As discussed in the previous section, the dispersion relations of phonons and phasons lead to Friedmann-like equations describing the expansion of the universe, contributing to the energy density and exhibiting different behaviors in various phases of cosmological evolution. In the spirit of [26], we propose that phonon and phason excitations can provide the seeds for composite matter formation. In this model, quasicrystal points act as point-particle excitations, forming composite particles such as hadrons. These excitations contribute to the radiation epoch spectrum, and as more composite matter emerges, the universe transitions to a matter-dominated epoch. As the universe expands, phonon excitations may continue to influence its evolution. In particular, the dispersion relation for phasons (Equation 12) includes an imaginary term that becomes significant at large wavelengths, leading to diffusive behavior, thereby excluding its contribution from the energy density.
3.1.1 Tiling Spaces: From Discrete to Continuous
While quantum field theory may be suitable for studying these transitions, most many-body systems, including quasicrystals and the universe, exist at finite temperatures. As a result, the equilibrium state of the system is described by a thermal density matrix, often involving mixed states. Consequently, conventional quantum field theory techniques may not be appropriate. The literature frequently employs effective field theory techniques to address these conditions. Here, we will describe tiling spaces using the probabilistic approach of [19]. Tiling spaces are defined by the continuous hull $X(\varLambda)$, which includes all translates of a point set $\varLambda$ in $\mathbb{R}^{d}$. This framework is typically applied to infinite quasicrystals, but to understand regimes where dynamics occur at scales much larger than the underlying spacing (potentially at the Planck scale), we expect to recover invariance under continuous translations in the tiling space.

### v2

@@ 2 The Universe Expansion from Quasicrystalline Tiling Spaces
The projection of the initial shells of the E8 lattice, when viewed from a 3D slice perspective, reveals distinct patterns that emerge through rescaling. In this process, the quasicrystal grows externally in size while simultaneously undergoing internal subdivision. See Figure (1) for a visual representation.
@@
The notion of tiling spaces emerges by treating each possible tiling configuration that can be constructed on a specific CPS as a point in this space. At each stage of the growth process, different tiling configurations are possible. Quantum mechanically, each possible tiling can be represented as a state, allowing us to describe the growth using the formalism of Hilbert spaces [17]. As we will show, tiling spaces offer a robust framework for capturing both the geometric and quantum mechanical properties of the universe’s expansion.
@@ 2.1 Universe as a Growing Quasicrystal Hypothesis
This model proposes an unconventional origin of the universe, diverging from the traditional Big Bang theory. It suggests that the universe began with the projection of a single point from a higher-dimensional lattice, such as the E8 root lattice. This initiated a sequence of point projections that sequentially constructed the universe’s structure. This process, governed by a computational framework, unfolds step by step transitioning from one shell $N_{\text{E8}}(n)$ to the next $N_{\text{E8}}(n+1)$ and so forth. The number of points in each shell grows rapidly, with the first ten shells containing 240, 2160, 6720, 17520, 30240, 60480, 82560, 140400, 181680, 272160 points, respectively. The projection into a 4D quasicrystalline structure not only expands the universe externally but also subdivides it internally. Points from newer shells intermingle with those from older shells, illustrating a complex process of growth and subdivision characteristic of the early universe (see Figure (2) for a lower-dimensional illustration and Figure (1) for 3D slices of the 4D ESQC).
Figure 2: A rotated Z2 is projected shell by shell with the color tracking the ordering. As this one‐dimensional universe expands, it also gets new points within the already projected region.
In this early stage, the universe consists of points, which are projections from the higher-dimensional root lattice. In the Lie algebra framework of modern gauge theories, these root lattice points represent fundamental particle states that form the Hilbert spaces of representations of the Lie algebra. At this stage, there is no scale or mass, and thus no distinction between particles, highlighting an analogy to the grand unification scale.
@@
Crystals and quasicrystals in condensed matter physics or solid state systems are known for their phonon quasiparticles, which emerge due to the breaking of translational symmetry. Quasicrystals, however, also host phasons—a less understood type of quasiparticle.
As the quasicrystalline universe evolves, phonons and phasons should emerge naturally due to the intrinsic properties of growing quasicrystals, even those composed of ’spacetime atoms.’ Phonons correspond to collective vibrational modes of the tiling configurations, while phasons are associated with the rearrangement of the quasicrystalline structure without changing the overall energy. The presence of these quasiparticles is significant, as their dispersion relations lead to a form of the Friedmann equation Equation (35), implying the establishment of standard expansion.
To incorporate phonons and phasons into our quantum description of tiling spaces, we can introduce creation and annihilation operators acting on the Hilbert space $\mathcal{H}_{n}$ at each growth step. Let $a^{\dagger}_{k,n}$ and $a_{k,n}$ denote the creation and annihilation operators for a phonon with wave vector $k$ at the $n$-th growth step, satisfying the commutation relations:
@@
2.2 Tiling Space
Let us switch now to a slight different formalist to demonstrate in a straightforward manner how energy density arises naturally from the underlying discrete structure in the regime of large number of points and tiles, which should be the regime where phonons and phasons emerge. We can start with the notion of tiling spaces using the probabilistic approach outlined in [14]. Tiling spaces are defined by the continuous hull $X(\varLambda)$, which includes all translations of a point set $\varLambda$ in $\mathbb{R}^{d}$. This framework is typically applied to infinite quasicrystals. However, to understand regimes where dynamics occur at scales much larger than the underlying spacing (potentially at the Planck scale), we expect to recover invariance under continuous translations in the tiling space.